# DSwinIR: Rethinking Window-based Attention for Image Restoration

### v2

@@ II-C Positioning Our Contribution
The prevailing trend in all-in-one image restoration has been to build sophisticated prompting superstructures on top of existing, often limited, backbones. Our work diverges from this trend by taking a first-principles approach. We posit that the next significant leap in performance should come not from more complex prompting schemes, but from a fundamental redesign of the attention mechanism at the heart of the backbone. We introduce the Deformable Sliding Window (DSwin) Attention, a new paradigm that is both token-centric and content-aware. It directly addresses the core architectural flaws of prior window-based models by eliminating boundary artifacts and enabling content-adaptive receptive fields. By focusing on evolving the foundational architecture, our DSwinIR demonstrates that a sufficiently intelligent and adaptive backbone can achieve state-of-the-art performance, even surpassing prompt-based methods. Our work does not preclude the use of prompts; rather, it provides a significantly more powerful and robust foundation upon which future, more effective prompting strategies can be built.
Figure 3: Overview of the proposed DSwinIR architecture. The model is built upon a U-shaped backbone where the core component is the DSwin Transformer Block (DSTB). The key modules include: (a) The Deformable Sliding Window Attention (DSwin) mechanism, which adaptively samples features by learning content-specific offsets. (b) The Multi-head Self-Attention (MS-DSA) module, which integrates multiple DSwin heads. (c) The Multi-Scale Gated Feed-Forward Network (MSG-FFN), which uses parallel convolutions to enhance feature extraction.
@@ III Method
In this section, we present our Deformable Sliding Window Transformer for Image Restoration (DSwinIR). We first provide an overview of the overall network architecture. Then, we elaborate on the heart of our model: the evolution from standard window attention to our proposed Deformable Sliding Window (DSwin) Attention, detailing its motivation and formulation. Finally, we describe the complete DSwin Transformer Block, which synergistically further extends the DSwin attention with a the practical multi-scale strategy.
@@ III-A Overall Architecture
Our architectural design is guided by a philosophy of proven effectiveness and hierarchical feature processing. To this end, we adopt the U-shaped encoder-decoder structure, as illustrated in Fig. 3. The core of this architecture is our novel DSwin Transformer Block, which we will detail in subsequent sections. The entire network is trained end-to-end, optimized with a simple yet robust $L_{1}$ pixel-wise loss function, which encourages sharp and accurate predictions by penalizing the absolute difference between the restored output $\hat{y}$ and the ground-truth image $y$:
$\mathcal{L}=||\hat{y}-y||_{1}.$ (1)
@@ III-B From Window Attention to Deformable Sliding Window Attention
The heart of our work lies in a reconceptualization of window-based attention. In the following subsections, we will introduce this principled evolution, starting from a critical analysis of standard window attention, then introducing our two key innovations—the token-centric sliding mechanism and the content-aware deformable sampling—which together form our final DSwin Attention.
@@ III-B1 Preliminaries: Revisiting Window-based Self-Attention
To fully appreciate the novelty of our approach, it is essential to first deconstruct the standard window-based self-attention mechanism, as depicted in Fig. 1(c). A query token near the center may enjoy a full set of neighbors, but one at the edge—and especially a corner—is starved of context, blind to potentially crucial information just across the arbitrary window boundary. This “boundary blindness” is the first fundamental flaw we aim to solve. We briefly formalize its operation here to precisely identify the source of its inherent architectural flaws.
@@
where $d$ is the head dimension and $\mathbf{B}$ is a learnable relative position bias. As discussed in the introduction, this formulation, while efficient, inherently suffers from boundary artifacts and relies on a content-agnostic, fixed receptive field. Consequently, two spatially adjacent tokens that happen to straddle a window boundary are rendered computationally invisible to each other, creating an artificial information bottleneck that prohibits direct interaction.
III-B2 A Token-Centric Paradigm: The Sliding Window Mechanism
The primary source of boundary artifacts is the window-centric partitioning scheme itself. Therefore, our first key innovation is to abandon this paradigm in favor of a token-centric approach, inspired by the fundamental sliding operation of convolution that has been central to computer vision for decades. Our goal is to ensure that every token in the feature map can serve as the center of its own attention computation, enjoying a complete, uninterrupted contextual neighborhood, much like the sliding operation of a convolution kernel.
[…]
@@ III-B2 A Token-Centric Paradigm: The Sliding Window Mechanism
where $(u,v)$ are the relative coordinates within the $k\times k$ neighborhood. In stark contrast to the segregated windowing scheme, the local neighborhoods of any two adjacent tokens, such as $\mathbf{q}_{i,j}$ and $\mathbf{q}_{i,j+1}$, now possess a substantial overlap. This shared context explicitly remedies the information bottleneck at artificial boundaries. This sliding mechanism thus naturally creates overlapping receptive fields, ensuring smooth and robust information propagation across all spatial locations.
III-B3 Content-Aware Sampling: The Deformable Mechanism
While the sliding window paradigm successfully resolves the boundary problem, its reliance on a rigid, rectilinear grid still represents a significant constraint. This fixed partitioning is fundamentally misaligned with the arbitrary geometry of natural image structures and degradation patterns, such as the fine, diagonal lines of rain streaks or the amorphous, non-uniform nature of haze. Such a mismatch compromises the model’s ability to efficiently aggregate contextually relevant features, pointing to a critical need for a more flexible mechanism that can dynamically shape its receptive field in a content-aware manner
@@ III-B3 Content-Aware Sampling: The Deformable Mechanism
The network $f_{\theta}$ is implemented as a depth-wise separable convolution followed by a linear projection, ensuring it remains computationally efficient. The deformed key and value features are then sampled from the original feature map at the new coordinates $(i+u+\Delta u_{i,j}^{(u,v)},j+v+\Delta v_{i,j}^{(u,v)})$. Since the offsets are fractional, we use differentiable bilinear interpolation to sample the features, ensuring the entire process is end-to-end trainable. A value at a fractional location $(x,y)$ is computed as:
@@
[…]
@@
where $\alpha_{i,j}^{(u,v)}$ represents the attention weights computed using the deformed keys, and $\mathbf{v}_{\text{interp}}$ denotes the value sampled via bilinear interpolation. This final formulation, as illustrated in Fig. 1(d), represents a holistic and synergistic solution. It is not merely a sequence of two independent improvements, but a unified paradigm where each component empowers the other. Thesliding mechanism lays a robust, boundary-free foundation, ensuring every token has equal access to its local context. Upon this foundation, the deformable mechanism acts with surgical precision, intelligently warping the sampling grid to focus on the most salient information, wherever it may lie. The result is a powerful and expressive attention mechanism that is both structurally sound and semantically aware—perfectly tailored for the complex demands of image restoration.
@@
Three-Task Restoration 3 Denoising, Dehazing, Deraining
BSD400 [75], WED [76], RESIDE (SOTS) [77], Rain100L [78], BSD68 [79]
Five-Task Restoration 5
Adds Motion Deblurring & Low-light
Enhancement to the three-task setup
Adds GoPro [80] and LOLv1 [81] to the three-task training set
Synthetic Deweathering 3 Snow, Raindrop, Outdoor Rain Snow100K [82], Raindrop [83], Outdoor-Rain [84]
Real-World Deweathering 3 Real Rain, Real Snow, Real Haze SPA-Data [17], RealSnow[85], REVIDE[86]
Composite Degradation 11
Low-light, Haze, Rain, Snow,
and their combinations
DIV2K [87] with corresponding degradations
@@ III-C DSwin Transformer Block
The complete DSwin Transformer Block is constructed by integrating our DSwin attention and feedforward network with a complementary multi-scale strategy, which provides much more robust feature extraction.
@@ III-C1 Multi-Scale DSwin (MS-DSwin) Attention
Image content and degradations manifest across a spectrum of scales. The standard multi-head attention framework provides a natural and computationally efficient vehicle for capturing such multi-scale information. Instead of treating all attention heads identically, we leverage this structure to explicitly assign different spatial extents to different groups of heads, as depicted in Fig. 3(b).
Specifically, for a model with $N$ heads, we partition them into groups, with each group $h$ being assigned a unique sliding window kernel size $k_{h}$. This allows the model to simultaneously aggregate features from both compact local neighborhoods and broader contextual regions within a single MS-DSwin module. The outputs of all heads are then concatenated and passed through a final linear projection, fusing the multi-scale information at negligible extra computational cost.
@@ III-C2 Multi-Scale Gated Feed-Forward Network (MSG-FFN)
A standard Feed-Forward Network (FFN) in a Vision Transformer processes each token’s features independently via point-wise linear layers, momentarily losing the rich spatial context captured by the attention layer. To counteract this information bottleneck and maintain a holistic multi-scale design throughout the entire block, we replace the standard FFN with our proposed Multi-Scale Gated FFN (MSG-FFN).
As shown in Fig. 3(c), after an initial channel expansion layer, the MSG-FFN splits the features into parallel branches. Each branch employs a depth-wise convolution with a different kernel size or dilation rate, explicitly re-introducing spatial reasoning at multiple scales. These parallel features are then passed through a gating mechanism and a final projection layer. This design ensures that both the attention and FFN components of our block actively engage in multi-scale spatial feature processing, creating a more powerful and synergistic architecture.
@@ III-D Remarks
It is crucial to position our contribution not as another incremental improvement, but as a fundamental shift in designing attention mechanisms for dense prediction tasks. Previous works often resort to heuristic patches like shifted windows or complex static patterns (e.g., axial stripes, cross-shapes) to mitigate the boundary artifacts of window attention. These methods, however, merely replace one content-agnostic pattern with another, more elaborate one. Our DSwin Attention, in contrast, is a first-principles solution. The token-centric sliding paradigm completely eliminates the concept of a boundary, and the deformable sampling makes the receptive field itself a function of the data. This addresses the core issues of boundary truncation and inflexible receptive fields directly and dynamically. DSwinIR achieves a more powerful synthesis of the core strengths of CNNs and Transformers.
The prevailing trend in all-in-one restoration has focused on building sophisticated prompting superstructures atop existing, often limited, backbones. The sliding window mechanism re-introduces a strong, locality-aware inductive bias, akin to convolution, which is highly effective for image processing. Simultaneously, the content-aware deformable sampling endows the model with powerful, data-driven adaptivity for long-range modeling, which is the hallmark of attention mechanisms. Our work demonstrates that evolving the foundational architecture itself is a more direct and potent path to state-of-the-art performance.
@@
TABLE II: Quantitative comparisons for Setting 1: three distinct degradation tasks. Results of our proposed DSwinIR are in bold. * denotes the results are adopted from existing work [15, 67].
@@
Type Method Venue Denoising (CBSD68[79]) Dehazing Deraining Average
$\sigma=15$ $\sigma=25$ $\sigma=50$ SOTS [77] Rain100L [78]
General
MPRNet [6] CVPR’21 33.27/0.920 30.76/0.871 27.29/0.761 28.00/0.958 33.86/0.958 30.63/0.894
Restormer [45] CVPR’22 33.72/0.930 30.67/0.865 27.63/0.792 27.78/0.958 33.78/0.958 30.75/0.901
NAFNet [34] ECCV’22 33.03/0.918 30.47/0.865 27.12/0.754 24.11/0.928 33.64/0.956 29.67/0.844
FSNet* [35] TPAMI’23 33.81/0.930 30.84/0.872 27.69/0.792 29.14/0.968 35.61/0.969 31.42/0.906
DRSformer* [13] CVPR’23 33.28/0.921 30.55/0.862 27.58/0.786 29.02/0.968 35.89/0.970 31.26/0.902
MambaIR* [50] ECCV’24 33.88/0.931 30.95/0.874 27.74/0.793 29.57/0.970 35.42/0.969 31.51/0.907
All-in-One
DL [88] TPAMI’19 33.05/0.914 30.41/0.861 26.90/0.740 26.92/0.391 32.62/0.931 29.98/0.875
AirNet [61] CVPR’22 33.92/0.932 31.26/0.888 28.00/0.797 27.94/0.962 34.90/0.967 31.20/0.910
IDR* [62] CVPR’23 33.89/0.931 31.32/0.884 28.04/0.798 29.87/0.970 36.03/0.971 31.83/0.911
PromptIR [15] NeurIPS’23 33.98/0.933 31.31/0.888 28.06/0.799 30.58/0.974 36.37/0.972 32.06/0.913
Gridformer* [89] IJCV’24 33.93/0.931 31.37/0.887 28.11/0.801 30.37/0.970 37.15/0.972 32.19/0.912
NDR [90] TIP’24 34.01/0.932 31.36/0.887 28.10/0.798 28.64/0.962 35.42/0.969 31.51/0.910
InstructIR [16] ECCV’24 34.15/0.933 31.52/0.890 28.30/0.804 30.22/0.959 37.98/0.978 32.43/0.913
TextualDegRemoval[66] CVPR’24 34.01/0.933 31.39/0.890 28.18/0.802 31.63/0.980 37.58/0.979 32.63/0.917
AdaIR [63] ICLR’25 34.12/0.935 31.45/0.892 28.19/0.802 31.06/0.980 38.64/0.983 32.69/0.918
Perceive-IR [67] TIP’25 34.13/0.934 31.53/0.890 28.31/0.804 30.87/0.975 38.29/0.980 32.63/0.917

### v3

@@ II-C Positioning Our Contribution
The power of content-adaptive receptive fields has been clearly established, first bringing significant gains to convolutional networks [DeformableConv_DaiQXLZHW17]. This principle was subsequently adapted for Transformer architectures, such as in the multi-scale deformable attention proposed by Zhu et al. [Zhu2021DeformableDETR] and the deformable attention mechanisms developed by Xia et al. [Xia2022DAT]. Drawing inspiration from this clear line of progress, we introduce the Deformable Sliding Window (DSwin) Attention, a new paradigm that is both token-centric and content-aware. It directly addresses the core architectural flaws of prior window-based models by eliminating boundary artifacts and enabling content-adaptive receptive fields. By focusing on evolving the foundational architecture, our DSwinIR demonstrates that a sufficiently intelligent and adaptive backbone can achieve state-of-the-art performance, even surpassing prompt-based methods. Our work does not preclude the use of prompts; rather, it provides a significantly more powerful and robust foundation upon which future, more effective prompting strategies can be built.
Figure 3: Overview of the proposed DSwinIR architecture. The model is built upon a U-shaped backbone where the core component is the DSwin Transformer Block (DSTB). The key modules include: (a) Deformable Sliding Window Attention (DSwin), which adaptively samples features by learning content-dependent offsets; (b) Multi-Scale Deformable Sliding Window (MS-DSwin) Attention, which integrates multi-scale DSwin attention across multiple attention heads; and (c) Multi-Scale Gated Feed-Forward Network (MSG-FFN), which leverages parallel convolutional branches to enhance feature representation.
@@ III Method
In this section, we introduce DSwinIR. We first outline the overall architecture, then describe the progression from standard window attention to the proposed DSwin attention, including the motivation and formulation. Finally, we present the DSwin Transformer Block and a practical multi‑scale design.
@@ III-A Overall Architecture
We adopt a U‑shaped encoder–decoder architecture with hierarchical feature processing (Fig. 3). The core of this architecture is our novel DSwin Transformer Block, which we will detail in subsequent sections. The entire network is trained end-to-end, optimized with the $L_{1}$ pixel-wise loss function, which encourages sharp and accurate predictions by penalizing the absolute difference between the restored output and the ground-truth image.
@@ III-B From Window Attention to Deformable Sliding Window Attention
We revisit window‑based attention and develop the Deformable Sliding Window (DSwin) attention from this perspective. In the following subsections, we outline the formulation and design choices.
@@ III-B1 Preliminaries: Revisiting Window-based Self-Attention
We first review standard window‑based self‑attention as depicted in Figure 1(c). A query token near the center may enjoy a full set of neighbors, but one at the edge—and especially a corner—is starved of context, blind to potentially crucial information just across the arbitrary window boundary. We refer to this as boundary blindness and aim to mitigate it. We briefly formalize its operation here to precisely identify the source of these inherent architectural flaws.
@@
where $d$ is the head dimension and $\mathbf{B}$ is a learnable relative position bias. As discussed in the introduction, this formulation, while efficient, inherently suffers from insufficient cross-window interaction.
III-B2 Token‑Centric Sliding Window
Boundary artifacts largely stem from non‑overlapping partitions. We replace window‑centric partitioning with a token‑centric sliding window, inspired by sliding convolution.
For a given query token at position $(i,j)$, we define its local neighborhood of size $k\times k$. This operation extracts a feature patch $\mathcal{N}_{k}(i,j)\in\mathbb{R}^{k\times k\times C}$ centered at each position $(i,j)$. It effectively transforms the feature map $\mathbf{X}$ into a sequence of patches, where each patch contains the keys and values for a corresponding query token. The attention for the query token $\mathbf{q}_{i,j}$ is then computed over the keys $\mathbf{K}_{i,j}$ and values $\mathbf{V}_{i,j}$ extracted from its neighborhood $\mathcal{N}_{k}(i,j)$:
@@ III-B2 Token‑Centric Sliding Window
where $(u,v)$ are the relative coordinates within the $k\times k$ neighborhood. Unlike non‑overlapping windows, this sliding mechanism reduces the information bottleneck at boundaries and facilitates communication across spatial locations.
III-B3 Content‑Aware Deformable Sampling
Sliding windows alleviate boundary issues but still rely on a fixed, rectangular sampling grid. This fixed partitioning is misaligned with the arbitrary geometry of natural image structures and degradation patterns, such as the fine, diagonal lines of rain streaks or the amorphous, non-uniform nature of haze. Such a mismatch compromises the model’s ability to efficiently aggregate contextually relevant features, pointing to a critical need for a more flexible mechanism that can dynamically shape its receptive field in a content-aware manner.
@@ III-B3 Content‑Aware Deformable Sampling
The network $f_{\theta}$ is implemented as a lightweight module consisting of a depth-wise separable convolution, followed by a GELU activation function, and a final linear projection. No explicit offset regularization was required in our experiments; offsets are learned end‑to‑end via the restoration loss. The deformed key and value features are then sampled from the original feature map at the new coordinates $(i+u+\Delta u_{i,j}^{(u,v)},j+v+\Delta v_{i,j}^{(u,v)})$. Since the offsets are fractional, we use differentiable bilinear interpolation to sample the features, ensuring the entire process is end-to-end trainable. A value at a fractional location $(x,y)$ is computed as:
@@
where $(i^{\prime},j^{\prime})\in\{(0,0),(0,1),(1,0),(1,1)\}$ are the four integer grid neighbors, and $w(i^{\prime},j^{\prime})$ are the bilinear interpolation weights, which are proportional to the distance from $(x,y)$ to the opposite corner.
@@
where $\alpha_{i,j}^{(u,v)}$ represents the attention weights computed using the deformed keys, and $\mathbf{v}_{\text{interp}}$ denotes the value sampled via bilinear interpolation. This formulation combines sliding and deformable sampling: sliding provides overlapping local context; deformable offsets adapt sampling to content, as illustrated in Figure 1(d).
Figure 4: Visualization of the content-aware sampling mechanism in DSwinIR. The figure presents a qualitative analysis of our proposed DSwin Attention. (a) The input image, with a red dot indicating the anchor location for local analysis. (b) The corresponding local sampling pattern, showing the reference grid, the adaptively sampled points, and the learned offsets. (c) The Deformation Magnitude Map, which visualizes the offset distance ($d=\sqrt{dx^{2}+dy^{2}}$) as a heatmap, where brighter intensity signifies a larger offset.
@@
Three-Task Restoration 3 Denoising, Dehazing, Deraining BSD400 [arbelaez2011contour], WED [ma2017waterloo], RESIDE (SOTS) [li2019benchmarking], Rain100L [yang2017deep], BSD68 [martin2001database1]
Five-Task Restoration 5 Adds Motion Deblurring & Low-Light Enhancement to the three-task setup Adds GoPro [nah2017deep] and LOLv1 [wei2018deep] to the three-task training set
Synthetic Deweathering 3 Snow, Raindrop, Outdoor Rain Snow100K [Snow100K_dataset], Raindrop [raindrop], Outdoor-Rain [test1_dataset_allweather]
Real-World Deweathering 3 Real Rain, Real Snow, Real Haze SPA-Data [SPANet_WangY0C0L19], RealSnow[weather_data], REVIDE[Zhang2021Realhaze]
Real-World Deweathering 3 All-Time Adverse Weather Types WeatherBench [Guan2025WeatherBench]
Composite Degradation 11 Low-Light, Haze, Rain, Snow, and their combinations DIV2K [Agustsson_2017_CVPR_Workshops] with corresponding degradations
@@ III-C DSwin Transformer Block
The DSwin Transformer Block integrates DSwin attention and a feed‑forward network with a complementary multi‑scale strategy to strengthen feature extraction.
@@ III-C1 Multi-Scale DSwin (MS-DSwin) Attention
Image content and degradations manifest across a spectrum of scales. The standard multi-head attention framework provides a natural and computationally efficient vehicle for capturing such multi-scale information. Instead of treating all attention heads identically, we leverage this structure to explicitly assign different spatial extents to different groups of heads, as depicted in Figure 3(b). Specifically, for a model with $N$ heads, we partition them into groups, with each group $h$ being assigned a unique sliding window kernel size $k_{h}$. This allows the model to simultaneously aggregate features from both compact local neighborhoods and broader contextual regions within a single MS-DSwin module. Outputs from all heads are concatenated and projected to fuse multi‑scale information at a small additional cost.
@@ III-C2 Multi-Scale Gated Feed-Forward Network (MSG-FFN)
A standard feed‑forward network (FFN) processes tokens independently via point‑wise linear layers, without explicit spatial context. To keep a multi‑scale design throughout the block, we introduce a Multi‑Scale Gated FFN (MSG‑FFN): after channel expansion, features split into parallel depth‑wise convolution branches with different kernel sizes or dilations, followed by gating and projection. As shown in Figure 3(c), after an initial channel expansion layer, the MSG-FFN splits the features into parallel branches. Each branch employs a depth-wise convolution with a different kernel size or dilation rate, explicitly re-introducing spatial reasoning at multiple scales.
@@ III-D Remarks
It is crucial to position our contribution not as another incremental improvement, but as a complementary perspective in designing attention mechanisms for dense prediction tasks. Previous works typically resort to hand-crafted patterns. Our DSwin Attention, in contrast, introduces the token-centric and content-aware paradigm. This addresses the core issues of cross-window interaction and inflexible receptive fields directly and adaptively.
Content-Aware Property: To provide intuition for the dynamic behavior of our proposed DSwin Attention, we present a visual analysis of its learned offsets in Figure 4. This visualization highlights two key properties of our module. First, at a micro-level (middle column), the learned offsets enable the sampling points to break free from the rigid reference grid, allowing the module to adaptively fetch features from more informative, non-adjacent locations. Second, at a macro-level (right column), the module exhibits a strong content-aware capability. We visualize the offset magnitude ($d=\sqrt{dx^{2}+dy^{2}}$) as a heatmap and observe a clear correlation with input complexity. The module learns significantly larger offsets for challenging regions (e.g., complex textures or heavy degradation), indicating an active search for relevant information over a wider area. Conversely, it reverts to a local operation (minimal offsets) in simple, uniform regions. This behavior validates that our DSwin Attention operates as a dynamic, content-aware operator rather than a static receptive field expander. We provide a more quantitative analysis of the resulting receptive field in our ablation study.
@@
TABLE II: Quantitative comparisons for Setting 1: three distinct degradation tasks. * denotes results reported by prior work [PromptIR, zhang2024perceive].
@@
Type Method Venue Denoising (CBSD68[martin2001database1]) Dehazing Deraining Average
$\sigma=15$ $\sigma=25$ $\sigma=50$ SOTS [li2019benchmarking] Rain100L [yang2017deep]
General MPRNet [zamir2021multi] CVPR’21 33.27/0.920 30.76/0.871 27.29/0.761 28.00/0.958 33.86/0.958 30.63/0.894
Uformer [Uformer] CVPR’22 33.55/0.922 30.71/0.859 27.63/0.786 27.13/0.953 34.28/0.961 30.66/0.896
Restormer [Restormer] CVPR’22 33.72/0.930 30.67/0.865 27.63/0.792 27.78/0.958 33.78/0.958 30.75/0.901
NAFNet [chen2022simple] ECCV’22 33.03/0.918 30.47/0.865 27.12/0.754 24.11/0.928 33.64/0.956 29.67/0.844
FSNet* [CuiRCK24] TPAMI’23 33.81/0.930 30.84/0.872 27.69/0.792 29.14/0.968 35.61/0.969 31.42/0.906
DRSformer* [DRSformer] CVPR’23 33.28/0.921 30.55/0.862 27.58/0.786 29.02/0.968 35.89/0.970 31.26/0.902
MambaIR* [MambaIR] ECCV’24 33.88/0.931 30.95/0.874 27.74/0.793 29.57/0.970 35.42/0.969 31.51/0.907
VmambaIR [Shi2025VmambaIR] TCSVT’25 33.75/0.927 31.10/0.879 27.81/0.782 29.75/0.970 35.63/0.970 31.61/0.906
MaIR [Li2025MaIR] CVPR’25 33.60/0.925 30.98/0.876 27.69/0.779 28.94/0.965 34.52/0.964 31.15/0.902
All-in-One DL [dl] TPAMI’19 33.05/0.914 30.41/0.861 26.90/0.740 26.92/0.391 32.62/0.931 29.98/0.875
AirNet [AirNet] CVPR’22 33.92/0.932 31.26/0.888 28.00/0.797 27.94/0.962 34.90/0.967 31.20/0.910
IDR* [IDR] CVPR’23 33.89/0.931 31.32/0.884 28.04/0.798 29.87/0.970 36.03/0.971 31.83/0.911
PromptIR [PromptIR] NeurIPS’23 33.98/0.933 31.31/0.888 28.06/0.799 30.58/0.974 36.37/0.972 32.06/0.913
Gridformer* [Gridformer] IJCV’24 33.93/0.931 31.37/0.887 28.11/0.801 30.37/0.970 37.15/0.972 32.19/0.912
NDR [NDR] TIP’24 34.01/0.932 31.36/0.887 28.10/0.798 28.64/0.962 35.42/0.969 31.51/0.910
InstructIR [InstructIR] ECCV’24 34.15/0.933 31.52/0.890 28.30/0.804 30.22/0.959 37.98/0.978 32.43/0.913
TextualDegRemoval[texture] CVPR’24 34.01/0.933 31.39/0.890 28.18/0.802 31.63/0.980 37.58/0.979 32.63/0.917
AdaIR [cui2025adair] ICLR’25 34.12/0.935 31.45/0.892 28.19/0.802 31.06/0.980 38.64/0.983 32.69/0.918
Perceive-IR [zhang2024perceive] TIP’25 34.13/0.934 31.53/0.890 28.31/0.804 30.87/0.975 38.29/0.980 32.63/0.917
PoolNet [PoolNet] TIP’25 34.10/0.935 31.45/0.892 28.18/0.803 30.94/0.980 38.54/0.983 32.64/0.919
VLU-Net [Zeng2025VLUNet] CVPR’25 34.13/0.935 31.48/0.892 28.23/0.804 30.71/0.980 38.93/0.984 32.70/0.919